OpenStudy (anonymous):

give the exact solutions and approximate solutions to three decimal points (-3)^2=1 i got (2,4) no approximates

OpenStudy (chaise):

Not sure what you're asking. Sorry.

OpenStudy (anonymous):

im asking if my answers are right to the problem i posted

OpenStudy (chaise):

Repost the question please

OpenStudy (anonymous):

(x-3)^2=1

OpenStudy (anonymous):

(x-3)^2=1 x² - 6x + 9 - 1 = 0 x² - 6x + 8 = 0 x² -4x -2x + 8 = 0 x(x-4) -2(x-4) = 0 (x-2)(x-4) = 0 so either x-2 = 0 or x-4 = 0 so x = 2 or x = 4 yes in this case u do not get approximations, maybe some of the questions might not have approximate roots that u can show in decimal numbers....